One positive number is one-fifth of another number. The difference between the two numbers is 136, find the numbers.
The two numbers in increasing order are _____ and _________

Answer by fractalier(6550)

Call the larger one x. Thus the smaller will be (1.5)x. Their difference is
x - (1/5)x = 136
(4/5)x = 136
x = 170
and then
(1/5)(170) = 34
In increasing order they are 34 and 170.